# SocialEyes: Scaling mobile eye-tracking to multi-person social settings

### v2

@@ Abstract.
$c$s: Human-centered computing Usability testingccs: Human-centered computing Collaborative interactionccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Visualization toolkitsccs: Human-centered computing Empirical studies in ubiquitous and mobile computingccs: Applied computing Performing artsccs: Software and its engineering Empirical software validation
@@ 1. Introduction
A schematic of our approach to solve the problem of multi-person eye-tracking in such collective viewing experiences is detailed in Figure 1. We imagine a scenario where multiple people observe a shared scene from different viewing angles. The eye movements of each viewer are recorded independently from their egocentric perspective (worldview) using commercial-grade eye-tracking glasses. However, to study collective gaze patterns, the wearer-specific gaze data needs to be synchronised and mapped onto a shared coordinate space (semantic gaze mapping). We record a centralview of the shared scene from a camera placed behind all observers. We employ Network Time Protocol (NTP) to temporally synchronise data from all devices. The synchronised gaze data from all viewers is then mapped to the shared coordinate space i.e. the centralview by finding robust feature matches between worldview–centralview image pairs and performing a projective transformation. Following our approach we design a software framework–SocialEyes. We implement and test SocialEyes during two public events in which we recorded eye data from 30 participants simultaneously. We detail our choices of eye-tracking hardware, the underlying infrastructure, and the implementation specifics of the framework. In validating our approach, we provide visualisation and analysis metrics for multi-person eye-tracking data. Overall, we make the following contributions in this paper:
@@
Figure 1. LEFT: Schematic of a shared scene depicting multiple people wearing mobile eye-tracking glasses gazing at a stage. A central camera at the back of the audience records the shared scene (centralview). MIDDLE: Egocentric worldviews of each glasses wearer need to be projected onto the centralview. Point correspondences between the egocentric and centralviews are used to calculate homography matrices that relate the two views. RIGHT: Using the computed homography matrices to remap the gaze coordinates, all individuals’ gaze points can be projected onto the shared scene (centralview).
@@ 3. SocialEyes: Design Concept
SocialEyes is a software framework that enables synchronous operation and analysis of multiple eye-tracking data streams in unrestricted, real-world settings. We followed the Research through Design approach (Stappers and Giaccardi, 2017) to design a software framework that is independent of specific hardware, network, and storage components, allowing it to adapt to various infrastructure constraints. In our software framework, presented in Figure 2, incoming data consists of three parallel streams: i) gaze coordinates from each eye-tracking device ii) worldview video recorded from a head-mounted camera on each device, and iii) a centralview video of the shared scene, captured and streamed by a stationary camera mounted to the centre of the scene. Incoming data streams are handled by separate modules that engage based on the mode of operation. Modules in our framework abstract information and allow isolated development and updates of separate components. Below, we describe the different modes of operation.
@@ 3.2.1. Time-synchronisation
. A critical challenge in managing multiple devices on a network is time synchronization, as each device relies on its internal clock to timestamp data samples. Independent internal clocks must be synchronised to a master clock to ensure accurate timestamp comparisons. All eye-tracking devices, recording cameras, and internal servers are, therefore, synchronised to a master NTP server accessible on the local network. This synchronisation ensures negligible clock offsets between devices, however, internal clocks on these devices can still accumulate variable drifts over long duration recordings. In the streaming mode of operation, this drift can be ignored for most applications to favour low-latency operation. However, for applications requiring fine-grained time alignment (typically in recording mode), we apply the clock-filter algorithm in NTP (Mills, 1994) to collect offset and drift at periodic intervals. Offsets between a reference device–the server running SocialEyes software– and each remote device (e.g., eye tracking glasses) are measured by exchanging a brief sequence of User Datagram Protocol (UDP) packets between the two involved devices. Each exchange returns a) an estimate of the round-trip-time (RTT) for the packet, and b) an estimate of the clock offset between the two devices at the time of the measurement, with the round-trip-time factored out. The final offset estimate is the one offset value for which the associated RTT value was minimal. Finally to align the timestamps from a device, its offset information is aggregated over the entire span of the recording (in recording mode) or a fixed size window (in streaming mode). An outlier-robust linear fit is calculated over the aggregated offset measurements using random sample consensus (Fischler and Bolles, 1981) to map the timestamps with the reference device. In our design, offset recording is performed as a background process by the GlassesRecord and GlassesStream modules for the recording and streaming mode respectively. The final synchronisation and linear mapping is performed as an optional process in each of the consumer modules i.e. Analysis, Visualisation, and Homography. We detail the functions of these modules below.
@@ 3.2.4. GlassesStream
GlassStream implements the streaming of gaze and worldview data streams in the streaming mode of operation. The module initialises a Kafka producer, parses incoming gaze and video data streams from selected eye-tracking devices on the network and publishes them to the Kafka cluster. The module also initiates a metrics server, incorporating Grafana and Prometheus services, to monitor the Kafka cluster.
@@ 3.2.5. Flask Monitor
The Flask Monitor module provides a real-time display of the incoming data streams to verify data integrity. The module initialises a Kafka consumer that reads each topic and updates corresponding graphs for that data stream. The module hosts a flask web-application that displays worldview streams from all devices with a configurable refresh rate (default 1Hz) and incoming frame rate (in FPS). Incoming gaze is displayed as a dynamically updating time-series of x and y coordinates, with a configurable refresh rate. The sample rates for gaze and worldview streams are set to a low value by default to reduce stress on the server and keep the monitor interface responsive. The incoming centralview is consumed in a separate process and displayed in a separate window, similar to the worldview frames. The module interface also features alert mechanisms to notify when data flow ceases or the frame rate drops.
@@ 3.2.6. Homography
The homography module implements semantic gaze mapping (as defined in our approach 1) from egocentric worldview coordinate space to centralview coordinate space simultaneously for multiple devices. Applying a pin-hole camera model, the worldview captured from each wearer’s perspective can be assumed to lie on a planar surface at an infinite distance. The projective transformation between the worldview and centralview for the same scene is then mathematically related by a planar homography. Incoming gaze samples (x, y coordinates) and worldview frames from each device are paired together with the nearest centralview frame using the marked timestamps from the respective input modules of each data stream. The module accommodates varying data sampling rates of the three data streams, since each stream’s sampling resolution depends upon factors like hardware configuration, recording parameters, and network bandwidth. The final temporal resolution of the homography module output is the same as the sampling rate of the provided centralview (i.e., for each centralview frame, the homography module attempts to map all gaze points from the provided worldviews+gaze pairs to that centralview frame). This mapping is accomplished by finding robust feature point correspondences between each worldview-centralview image pair. Matched key points are then used to calculate a homography matrix that is finally used to transform gaze from worldview to centralview coordinate space.
@@ 3.2.7. Visualisation
The visualisation module aligns timestamped worldview frames, gaze samples, and centralview frames with an additional input of the transformed gaze (in the centralview space) computed by the homography module. We design the following visualisation modes in this module:
  • Single-person gaze and worldview: Overlays gaze synchronised to the worldview stream for each selected viewer as a separate video file with an optional overlay of associated timestamps for each data sample.
  • Multi-person gaze and worldview: Renders a single video with all selected viewers’ worldviews with the corresponding gaze points overlay (as in the previous option) in a 2D grid of customisable size.
  • Single-person egocentric and transformed gaze view: Stacks the time synchronised worldview and centralview frames for a single viewer along with the corresponding gaze and transformed gaze overlay on each view respectively.
@@
  • Multi-person egocentric views with transformed gaze on centralview: Combines the worldviews from all selected viewers and plots it as a grid with the time synchronised centralview in the centre cell. Gaze for each viewer is overlaid in their respective worldview and the centre cell displays all the transformed gaze points similar to the previous option.
@@ 4.1.1. Apparatus
  • Eye-tracking equipment: We employed Pupil Labs Neon eye-tracking devices (Baumann and Dierkes, 2023) that apply a deep learning pipeline to provide reliable gaze predictions in unrestricted settings, requiring minimal calibration and setup. At the core of these eye trackers is a hardware module equipped with binocular infrared cameras that capture eye images with a resolution of 192 x 192 pixels at 200 Hz and a front-facing RGB camera that records the worldview with a resolution of 1600 x 1200 pixels at 30 Hz. The module is mounted to a 3D-printed, wearable frame that connects to an accompanying Android smartphone (Motorola Edge 40 Pro). The minimal form factor of the frames is designed to be unobtrusive and effortless for participants, making them a good fit for longer-duration studies. The accompanying Android phone runs a proprietary vendor application–“Neon Companion”–that hosts a suite of HTTP REST Application Programming Interfaces (APIs) enabling remote streaming of data and controlling operations such as start/stop recording. During the data collection event, each pair of eye-tracking glasses was connected to the companion smartphone device running Pupil’s Labs NEON Companion android application in the background. The phones were locked and connected to an Anker 6-in-1 USB-C hub that extended the ports to connect them to ethernet, power, and the NEON glasses concurrently. Each smartphone was secured in a bag attached to the respective participant’s chair, ensuring free movement of the seated participant (see Figure 3B). The ethernet cable connected the device to a network switch placed under the false floor, in the centre of the audience rows. The global NTP server for all smartphones, the central camera, and the server computers were all set to the University’s NTP server accessible on the local network.
@@
  • Network Infrastructure: Each pair of Neon glasses requires a bandwidth of 6-8 Megabytes per second (MBps) to stream gaze and worldview data. For real-time streaming operations, it is critical to consider the available bandwidth on a network to identify how many concurrent eye-tracking glasses’ data can be streamed. In this test, we used the recording mode of operation which has lower bandwidth requirements, since real-time data streams do not need to be accessed on the server. Nonetheless, we leveraged the stationary positioning of participants in our setup to provide precise timing information and optimal network reliability by connecting each eye-tracking device to a network switch (TP Link TL-SG1048) via ethernet. The network did not have internet access, to disable any background network activity on the connected Android smartphones.
  • Storage: Worldview videos and gaze data were recorded and stored locally on each smartphone through the Neon companion app. The recorded data was later uploaded to Pupil Cloud for post-processing of the recorded raw data and then downloaded for further analysis. Time offsets and logs from the GlassesRecord module were stored locally on the server running the module. Centralview video and metrics recorded by the CentralCam module were stored locally on the server running the CentralCam module.
@@ 4.1.2. SocialEyes implementation in Recording mode
  • GlassesRecord: We implemented the GlassesRecord module utilising the pupil-labs-real-time-api python library (https://pupil-labs-realtime-api.readthedocs.io) to send start, stop, discard, and save recording-triggers to each Neon device. Periodic clock offset measurements were recorded every ten seconds. Additionally, since the accompanying smartphones use Android as operating system, we used Android Debug Bridge (ADB) to retrieve status metrics including battery level, storage level, network ping, ADB server status, and connected USB devices (see Figure 4A). The status metrics facilitated continuous monitoring of devices and recording status. ADB was also used to remotely control the Android smartphones for remote troubleshooting and verifying indicators of failed or corrupted recordings that were not reported by the vendor application. These operations were important to prevent data loss and ensure that we would never need to interrupt the live social event to access a participants’ phone in the event of an error. The operator interface for this module was implemented as a Terminal User Interface (TUI) built using the python Textual library (https://textual.textualize.io/), featuring low runtime requirements and cross platform operation. The TUI (see Figure 4) displyed a list of all eye-tracking devices on the local network, along with accompanying colour-coded status fields for each device. An action panel at the bottom displayed a list of operations (actions) that can be performed on selected devices by pressing the corresponding action key. The TUI allowed selecting multiple devices to perform an action, including hotkey actions of Spacebar to select a single device; Shift + Arrow Up/Down: to expand selection up/down; Ctrl+A to select all devices; and Exclamation mark (!) to reverse selection.
    Figure 4. Data recording and processing interfaces. A Terminal User Interface (TUI) for the GlassesRecord module used to monitor and operate all 30 devices concurrently. B Command Line Interface to operate offline analysis and visualisation of the recorded data.
[…]
    Interfaces
@@
  • Homography, Analysis, and Visualisation: We applied a pre-trained Graph Neural Network–-SuperPoint/ SuperGlue (Sarlin et al., 2020)–to find robust feature point correspondences in our Homography module. Blinks were identified using the Neon blink-detection algorithm and filtered out of the raw gaze data before homography, analysis and visualisation. The operator interface for these modules was implemented as a command-line utility using the Python questionary (questionary.readthedocs.io) library (see Figure 4B). The command-line utility offered formatting of the directory structure for compatibility with modules functions, selecting a subset of recordings to perform functions, such as homography, visualisation, gaze metrics calculation, etc., and storing output to the local filesystem.
  • Visual angle calculation: Visual angle conversion was done by estimating the distances of each of the three participant rows from the videowall. The distances were 758 cm for the first, 885 cm for the second, and 1012 cm for the third participant row. Since the physical dimensions of the video wall were 384.5 cm for width and 216 cm for height, the diagonal (441.02 cm) projected a visual angle of 32.4° for the first, 28° for the second, and 25.6° for the third row. From the worldview recordings of participants in each row, the diagonal of the videowall was estimated to be 515 pixels for the first, 441 pixels for the second, and 396 pixels for the third row. The visual angle to pixel ratio was therefore calculated to be approximately 16 (i.e. one visual degree was represented by roughly 16 pixels in the 1600x1200 resolution worldview recordings).

### v3

@@ Abstract.
ccs: Human-centered computing Usability testingccs: Human-centered computing Collaborative interactionccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Visualization toolkitsccs: Human-centered computing Empirical studies in ubiquitous and mobile computingccs: Applied computing Performing artsccs: Software and its engineering Empirical software validation
@@ 1. Introduction
A schematic of our approach to solve the problem of multi-person eye-tracking in such collective viewing experiences is detailed in Figure 1. We imagine a scenario where multiple people observe a shared scene from different viewing angles. The eye movements of each viewer are recorded independently from their egocentric perspective (egoview) using commercial-grade eye-tracking glasses. However, to study collective gaze patterns, the wearer-specific gaze data needs to be synchronised and mapped onto a shared coordinate space (semantic gaze mapping). We record a centralview of the shared scene from a camera placed behind all observers. We employ Network Time Protocol (NTP) to temporally synchronise data from all devices. The synchronised gaze data from all viewers is then mapped to the shared coordinate space i.e. the centralview by finding robust feature matches between egoview–centralview image pairs and performing a projective transformation. Following our approach we design a software framework–SocialEyes. We implement and test SocialEyes during two public events in which we recorded eye data from 30 participants simultaneously. We detail our choices of eye-tracking hardware, the underlying infrastructure, and the implementation specifics of the framework. In validating our approach, we provide visualisation and analysis metrics for multi-person eye-tracking data. Overall, we make the following contributions in this paper:
@@
Figure 1. LEFT: Schematic of a shared scene depicting multiple people wearing mobile eye-tracking glasses gazing at a stage. A central camera at the back of the audience records the shared scene (centralview). MIDDLE: Egoviews of each glasses wearer need to be projected onto the centralview. Point correspondences between the egocentric and centralviews are used to calculate homography matrices that relate the two views. RIGHT: Using the computed homography matrices to remap the gaze coordinates, all individuals’ gaze points can be projected onto the shared scene (centralview).
@@ 3. SocialEyes: Design Concept
SocialEyes is a software framework that enables synchronous operation and analysis of multiple eye-tracking data streams in unrestricted, real-world settings. We followed the Research through Design approach (Stappers and Giaccardi, 2017) to design a software framework that is independent of specific hardware, network, and storage components, allowing it to adapt to various infrastructure constraints. In our software framework, presented in Figure 2, incoming data consists of three parallel streams: i) gaze coordinates from each eye-tracking device ii) egoview video recorded from a head-mounted camera on each device, and iii) a centralview video of the shared scene, captured and streamed by a stationary camera mounted to the centre of the scene. Incoming data streams are handled by separate modules that engage based on the mode of operation. Modules in our framework abstract information and allow isolated development and updates of separate components. Below, we describe the different modes of operation.
@@ 3.2.1. Time-synchronisation
A critical challenge in managing multiple devices on a network is time synchronization, as each device relies on its internal clock to timestamp data samples. Independent internal clocks must be synchronised to a master clock to ensure accurate timestamp comparisons. All eye-tracking devices, recording cameras, and internal servers are, therefore, synchronised to a master NTP server accessible on the local network. This synchronisation ensures negligible clock offsets between devices, however, internal clocks on these devices can still accumulate variable drifts over long duration recordings. In the streaming mode of operation, this drift can be ignored for most applications to favour low-latency operation. However, for applications requiring fine-grained time alignment (typically in recording mode), we apply the clock-filter algorithm in NTP (Mills, 1994) to collect offset and drift at periodic intervals. Offsets between a reference device–the server running SocialEyes software– and each remote device (e.g., eye tracking glasses) are measured by exchanging a brief sequence of User Datagram Protocol (UDP) packets between the two involved devices. Each exchange returns a) an estimate of the round-trip-time (RTT) for the packet, and b) an estimate of the clock offset between the two devices at the time of the measurement, with the round-trip-time factored out. The final offset estimate is the one offset value for which the associated RTT value was minimal. Finally to align the timestamps from a device, its offset information is aggregated over the entire span of the recording (in recording mode) or a fixed size window (in streaming mode). An outlier-robust linear fit is calculated over the aggregated offset measurements using random sample consensus (Fischler and Bolles, 1981) to map the timestamps with the reference device. In our design, offset recording is performed as a background process by the GlassesRecord and GlassesStream modules for the recording and streaming mode respectively. The final synchronisation and linear mapping is performed as an optional process in each of the consumer modules i.e. Analysis, Visualisation, and Homography. We detail the functions of these modules below.
@@ 3.2.4. GlassesStream
GlassStream implements the streaming of gaze and egoview data streams in the streaming mode of operation. The module initialises a Kafka producer, parses incoming gaze and video data streams from selected eye-tracking devices on the network and publishes them to the Kafka cluster. The module also initiates a metrics server, incorporating Grafana and Prometheus services, to monitor the Kafka cluster.
@@ 3.2.5. Flask Monitor
The Flask Monitor module provides a real-time display of the incoming data streams to verify data integrity. The module initialises a Kafka consumer that reads each topic and updates corresponding graphs for that data stream. The module hosts a flask web-application that displays egoview streams from all devices with a configurable refresh rate (default 1Hz) and incoming frame rate (in FPS). Incoming gaze is displayed as a dynamically updating time-series of x and y coordinates, with a configurable refresh rate. The sample rates for gaze and egoview streams are set to a low value by default to reduce stress on the server and keep the monitor interface responsive. The incoming centralview is consumed in a separate process and displayed in a separate window, similar to the egoview frames. The module interface also features alert mechanisms to notify when data flow ceases or the frame rate drops.
@@ 3.2.6. Homography
The homography module implements semantic gaze mapping (as defined in our approach 1) from egoview coordinate space to centralview coordinate space simultaneously for multiple devices. Applying a pin-hole camera model, the egoview captured from each wearer’s perspective can be assumed to lie on a planar surface at an infinite distance. The projective transformation between the egoview and centralview for the same scene is then mathematically related by a planar homography. Incoming gaze samples (x, y coordinates) and egoview frames from each device are paired together with the nearest centralview frame using the marked timestamps from the respective input modules of each data stream. The module accommodates varying data sampling rates of the three data streams, since each stream’s sampling resolution depends upon factors like hardware configuration, recording parameters, and network bandwidth. The final temporal resolution of the homography module output is the same as the sampling rate of the provided centralview (i.e., for each centralview frame, the homography module attempts to map all gaze points from the provided egoviews+gaze pairs to that centralview frame). This mapping is accomplished by finding robust feature point correspondences between each egoview-centralview image pair. Matched key points are then used to calculate a homography matrix that is finally used to transform gaze from egoview to centralview coordinate space.
@@ 3.2.7. Visualisation
The visualisation module aligns timestamped egoview frames, gaze samples, and centralview frames with an additional input of the transformed gaze (in the centralview space) computed by the homography module. We design the following visualisation modes in this module:
  • Single-person gaze and egoview: Overlays gaze synchronised to the egoview stream for each selected viewer as a separate video file with an optional overlay of associated timestamps for each data sample.
  • Multi-person gaze and egoview: Renders a single video with all selected viewers’ egoviews with the corresponding gaze points overlay (as in the previous option) in a 2D grid of customisable size.
  • Single-person egocentric and transformed gaze view: Stacks the time synchronised egoview and centralview frames for a single viewer along with the corresponding gaze and transformed gaze overlay on each view respectively.
@@
  • Multi-person egocentric views with transformed gaze on centralview: Combines the egoviews from all selected viewers and plots it as a grid with the time synchronised centralview in the centre cell. Gaze for each viewer is overlaid in their respective egoview and the centre cell displays all the transformed gaze points similar to the previous option.
@@ 4.1.1. Apparatus
  • Eye-tracking equipment: We used the Pupil Labs Neon eye-tracking devices (Baumann and Dierkes, 2023) that apply a deep learning pipeline to provide reliable gaze predictions in unrestricted settings. At the core of these eye trackers is a hardware module equipped with binocular infrared cameras that capture eye images with a resolution of 192 x 192 pixels at 200 Hz and a front-facing RGB camera that records the egoview with a resolution of 1600 x 1200 pixels at 30 Hz. The fixed geometry of eye and egoview cameras on the module eliminates the requirement of explicit eye-egoview calibration before each recording; the gaze is recorded in egoview coordinates. The module is mounted to a 3D-printed, wearable frame that connects to an accompanying Android smartphone (Motorola Edge 40 Pro). The minimal form factor of the frames is designed to be unobtrusive and effortless for participants, making them a good fit for longer-duration studies. The accompanying Android phone runs a proprietary vendor application–“Neon Companion”–that hosts a suite of HTTP REST Application Programming Interfaces (APIs) enabling remote streaming of data and controlling operations such as start/stop recording. During the data collection event, each pair of eye-tracking glasses was connected to the companion smartphone device running Pupil’s Labs NEON Companion android application in the background. The phones were locked and connected to an Anker 6-in-1 USB-C hub that extended the ports to connect them to ethernet, power, and the NEON glasses concurrently. Each smartphone was secured in a bag attached to the respective participant’s chair, ensuring free movement of the seated participant (see Figure 3B). The ethernet cable connected the device to a network switch placed under the false floor, in the centre of the audience rows. The global NTP server for all smartphones, the central camera, and the server computers were all set to the University’s NTP server accessible on the local network.
@@
  • Network Infrastructure: Each pair of Neon glasses requires a bandwidth of 6-8 Megabytes per second (MBps) to stream gaze and egoview data. For real-time streaming operations, it is critical to consider the available bandwidth on a network to identify how many concurrent eye-tracking glasses’ data can be streamed. In this test, we used the recording mode of operation which has lower bandwidth requirements, since real-time data streams do not need to be accessed on the server. Nonetheless, we leveraged the stationary positioning of participants in our setup to provide precise timing information and optimal network reliability by connecting each eye-tracking device to a network switch (TP Link TL-SG1048) via ethernet. The network did not have internet access, to disable any background network activity on the connected Android smartphones.
  • Storage: egoview videos and gaze data were recorded and stored locally on each smartphone through the Neon companion app. The recorded data was later uploaded to Pupil Cloud for post-processing of the recorded raw data and then downloaded for further analysis. Time offsets and logs from the GlassesRecord module were stored locally on the server running the module. Centralview video and metrics recorded by the CentralCam module were stored locally on the server running the CentralCam module.
@@ 4.1.2. SocialEyes implementation in Recording mode
  • GlassesRecord: We implemented the GlassesRecord module utilising the pupil-labs-real-time-api python library (https://pupil-labs-realtime-api.readthedocs.io) to send start, stop, discard, and save recording-triggers to each Neon device. Periodic clock offset measurements were recorded every ten seconds. Additionally, since the accompanying smartphones use Android as operating system, we used Android Debug Bridge (ADB) to retrieve status metrics including battery level, storage level, network ping, ADB server status, and connected USB devices. The status metrics facilitated continuous monitoring of devices and recording status. ADB was also used to remotely control the Android smartphones for remote troubleshooting and verifying indicators of failed or corrupted recordings that were not reported by the vendor application. These operations were important to prevent data loss and ensure that we would never need to interrupt the live social event to access a participants’ phone in the event of an error. The operator interface for this module was implemented as a Terminal User Interface (TUI) built using the python Textual library (https://textual.textualize.io/), providing low runtime requirements and cross platform operation. The TUI displays a list of all eye-tracking devices on the local network, along with colour-coded status metrics for each device (battery/storage level, network ping, ADB server status, recording status, etc.). Multiple devices can be selected from this list to perform a desirable action with dedicated keystrokes. The following actions can be performed on all selected devices using the TUI: Start Recording; Save Recording (end and save recordings locally); Cancel Recording (end and discard ongoing recordings); Restart App (use ADB to restart the Neon Companion app, critical for troubleshooting app crashes remotely); Reconnect ADB (reconnect to wireless ADB, critical for recovering lost connection to devices).
@@
  • Homography, Analysis, and Visualisation: We applied a pre-trained Graph Neural Network–-SuperPoint/ SuperGlue (Sarlin et al., 2020)–to find robust feature point correspondences in our Homography module. Blinks were identified using the Neon blink-detection algorithm (Labs, 2023) and filtered out of the raw gaze data before homography, analysis and visualisation. The operator interface for these modules was implemented as a command-line utility using the Python questionary (questionary.readthedocs.io) library. The command-line utility offered formatting of the directory structure for compatibility with modules functions, selecting a subset of recordings to perform functions, such as homography, visualisation, gaze metrics calc‘ulation, etc., and storing output to the local filesystem. Gaze heatmaps were generated using the viridis scientific colormap (van der Walt et al., 2020), which is designed for accurate perceptual uniformity and accessibility for viewers with color vision deficiencies.
  • Visual angle calculation: Visual angle conversion was done by estimating the distances of each of the three participant rows from the videowall. The distances were 758 cm for the first, 885 cm for the second, and 1012 cm for the third participant row. Since the physical dimensions of the video wall were 384.5 cm for width and 216 cm for height, the diagonal (441.02 cm) projected a visual angle of 32.4° for the first, 28° for the second, and 25.6° for the third row. From the egoview recordings of participants in each row, the diagonal of the videowall was estimated to be 515 pixels for the first, 441 pixels for the second, and 396 pixels for the third row. The visual angle to pixel ratio was therefore calculated to be approximately 16 (i.e. one visual degree was represented by roughly 16 pixels in the 1600x1200 resolution egoview recordings).